An antacid tablet contains Magnesium hydroxide and an inert binding agent. The mass of the Magnesium hydroxide in the tablet was determined by titration with HCl. If one tablet required 39.1 mL of 0.2056 N HCl for neutralization; how many equivalents of the base are in the tablet ? How many grams is this ?



1
Expert's answer
2021-10-29T02:20:41-0400

Complete neutralization means that equivalents of a base and an acid are equal. Therefore,

"eq(Mg(OH)_2)=eq(HCl)=0.0391\ L\times\frac{0.2056\ eq(HCl)}{1\ L}=0.00804\ eq"

Since Mg(OH)2 is a dibasic base, one mole of it corresponds to two equivalents. The molar mass of it equals 58.32 g/mol. Therefore,

"m(Mg(OH)_2)=0.00804\ eq(Mg(OH)_2)\times\frac{1\ mol(Mg(OH)_2)}{2\ eq(Mg(OH)_2)}\times\frac{58.32\ g(Mg(OH)_2)}{1\ mol(Mg(OH)_2)}=0.234\ g"


Answer: 0.00804 eq; 0.234 g